# The direct linearization scheme with the Lamé function: The KP equation and reductions

### v2

@@ B.1 Distinct real cube roots
(1) In fact, in case equation (B.1) having distinct real roots within the real period $(0,2w_{1}]$, ′′$\wp^{\prime\prime}(z)$ will have at least one real zero. However, recalling ′′$\wp^{\prime\prime}(z)=6\wp^{2}(z)-\frac{g_{2}}{2}$ and $\wp^{2}(z)\geq e_{1}^{2}>\frac{g_{2}}{12}$ when $\Delta>0$ and $z\in(0,2w_{1})$ (see Proposition A.2), we know that ′′$\wp^{\prime\prime}(z)>0$. This means there is no zeros for ′′$\wp^{\prime\prime}(z)$ in $(0,2w_{1})$.
(2) In the case $\Delta<0$, to ensure that ′′$\wp^{\prime\prime}(z)=0$ has real roots, we only consider the case where $g_{2}>0$. When $g_{3}\geq 0$, according to the Proposition A.2, $\wp^{2}(z)\geq e_{3}^{2}>\frac{g_{2}}{12}$ always holds true within the real period. For the case $g_{3}<0$, from the equation (A.5), it is known that $\tilde{e}_{3}<-\sqrt{\frac{g_{2}}{12}}$. Due to $\lim_{z\to 0^{+}}\wp(z)=\lim_{z\to 0^{-}}\wp(z)=+\infty$, it can be deduced that $\wp(z)=\pm\sqrt{\frac{g_{2}}{12}}$ has four distinct real roots within the real period $(-\tilde{w}_{3},\tilde{w}_{3})$. In other words, $\wp^{\prime}(z)$ has four extremum points, which can be denoted as $-z_{1},-z_{2},z_{2},z_{1}$ in ascending order within the real period. Since $\lim_{z\to 0^{-}}\wp^{\prime}(z)=+\infty$, $\lim_{z\to 0^{+}}\wp^{\prime}(z)=-\infty$ and $\wp^{\prime}(\tilde{w}_{3})=\wp^{\prime}(-\tilde{w}_{3})=0$, it can be deduced that by selecting the appropriate $k$ such that $-\wp^{\prime}(z_{2})<|\wp^{\prime}(k)|<-\wp^{\prime}(z_{1})$ holds, and then equation (B.1) can have three distinct real roots.
@@
Meanwhile, as $x_{0}\to 0^{+}$, $g^{\prime}(x_{0})<0$. Therefore, when we take $-2\wp^{\prime\prime\prime}(\frac{w_{1}}{2})/\wp^{\prime\prime}(\frac{w_{1}}{2})\leq b_{0}\ll\infty$, we can arrive at the numerical values of the two extremum points of $g(x)$. By ensuring that $g(l)$ is situated between the local maximum and minimum values, we can obtain three zeros $\omega_{0}(l)=l$, $\omega_{1}(l)$ and $\omega_{2}(l)$ of equation (3.36). The numerical results can be calculated using Mathematica.

### v3

@@ B.1 Distinct real cube roots
(1) In fact, in case equation (B.1) having distinct real roots within the real period $(0,2w_{1}]$, $\wp^{\prime\prime}(z)$ will have at least one real zero. However, recalling $\wp^{\prime\prime}(z)=6\wp^{2}(z)-\frac{g_{2}}{2}$ and $\wp^{2}(z)\geq e_{1}^{2}>\frac{g_{2}}{12}$ when $\Delta>0$ and $z\in(0,2w_{1})$ (see Proposition A.2), we know that $\wp^{\prime\prime}(z)>0$. This means there is no zeros for $\wp^{\prime\prime}(z)$ in $(0,2w_{1})$.
(2) In the case $\Delta<0$, to ensure that $\wp^{\prime\prime}(z)=0$ has real roots, we only consider the case where $g_{2}>0$. When $g_{3}\geq 0$, according to the Proposition A.2, $\wp^{2}(z)\geq e_{2}^{2}>\frac{g_{2}}{12}$ always holds true within the real period. For the case $g_{3}<0$, from the equation (A.5), it is known that $\tilde{e}_{2}<-\sqrt{\frac{g_{2}}{12}}$. Due to $\lim_{z\to 0^{+}}\wp(z)=\lim_{z\to 0^{-}}\wp(z)=+\infty$, it can be deduced that $\wp(z)=\pm\sqrt{\frac{g_{2}}{12}}$ has four distinct real roots within the real period $(-\tilde{w}_{2},\tilde{w}_{2})$. In other words, $\wp^{\prime}(z)$ has four extremum points, which can be denoted as $z_{1},-z_{2},z_{2},z_{1}$ in ascending order within the real period. Since $\lim_{z\to 0^{-}}\wp^{\prime}(z)=+\infty$, $\lim_{z\to 0^{+}}\wp^{\prime}(z)=$-$\infty$ and $\wp^{\prime}(\tilde{w}_{2})=\wp^{\prime}(-\tilde{w}_{2})=0$, it can be deduced that by selecting the appropriate $k$ such that $\wp^{\prime}(z_{2})<|\wp^{\prime}(k)|<-\wp^{\prime}(z_{1})$ holds, and then equation (B.1) can have three distinct real roots.
@@
Meanwhile, as $x_{0}\to 0^{+}$, $g^{\prime}(x_{0})<0$. Therefore, when we take $2\wp^{\prime\prime\prime}(\frac{w_{1}}{2})/\wp^{\prime\prime}(\frac{w_{1}}{2})\leq b_{0}\ll\infty$, we can arrive at the numerical values of the two extremum points of $g(x)$. By ensuring that $g(l)$ is situated between the local maximum and minimum values, we can obtain three zeros $\omega_{0}(l)=l$, $\omega_{1}(l)$ and $\omega_{2}(l)$ of equation (3.36). The numerical results can be calculated using Mathematica.
Appendix C On elliptic solitons
The authors are grateful to the referee for the comments on elliptic solutions and related references, which inspire us to complete this brief review on elliptic solitons.
The investigation of “elliptic solitons” dates back to the research of Krichever [28] on elliptic solutions of the KP equation and the KdV equation. He proposed a criterion for such solutions that the solutions should be elliptic in the spatial variable $x$. For the KP equation, Krichever proved its elliptic solutions should take the form ([28], Theorem 4)
$u=\mathrm{const.}+\sum_{j=1}^{n}\wp(x-x_{j}(y,t)),$ (C.1)
where $\{x_{j}(y,t)\}$ obey some dynamical system of finite dimension. He also showed that (C.1) can be expressed in terms of $\theta$-function ([28], Corollary 3). The reduction to the KdV equation recovers the 2-gap elliptic solution in [10] and the results of the paper [4] that for the first time links the elliptic solutions of integrable equations with the theory of finite-dimensional dynamic systems.
The term “elliptic soliton” is introduced by Verdier in [64], where he described solitons using the language of algebraic geometry. In his definition, “algebraic solitons” are governed by some linear algebraic subgroup $G$, while when $G$ is an elliptic curve, the solutions are called “elliptic solitons”. Verdier pointed out in [64] that the elliptic solitons of the KP equation must be of the form
$u=\mathrm{const.}+2\sum_{j=1}^{n}m_{j}\wp(x-x_{j}(y,t)),$ (C.2)
where $m_{j}$ are triangular numbers, and some new elliptic source potentials (without time evolution) for the KdV equation were listed out in [64]. Verdier’s research was elaborated later with his collaborator Treibich in [58, 62, 63, 59] by using the tool of “tangential cover”. In particular, in [63] they presented the following elliptic source potential for the KdV equation:
$u=\sum_{j=0}^{3}a_{j}(a_{j}+1)\wp(x-x_{j}),~~~~~\sum_{j=0}^{3}a_{j}(a_{j}+1)=2n,$ (C.3)
where $n$ is the degree of the tangential cover. This is more general than the Lamé-Ince potential [26] which reads $u=g(g+1)\wp(x),~g\in\mathbb{N}$. The function (C.3) is known as the Treibich-Verdier potential and now the Darboux-Treibich-Verdier (DTV) potential. It is Matveev and Smirnov [35] who pointed out that Darboux found the same potential in 1882 [9] but given in terms of Jacobi elliptic functions. They also showed in [35] that the Darboux equation is a canonical form the de Sparre equation [8]. There are other mathematicians and physicists joined the research of the DTV potential, e.g. [6, 20, 50, 51, 53, 57, 65]. Later, finite-gap elliptical potentials of a more general form (e.g. [54], Eq.(0.10)) of the Schrödinger spectral problem $\psi_{xx}(x,\lambda)+u(x)(x,\lambda)=\lambda\psi(x,\lambda)$ were found [54, 61]. It is also remarkable that Gesztesy and Weikard [21] investigated the Schrödinger spectral problem by utilizing a theorem of Picard which concerned with the existence of solutions that are elliptic of the $n$th-order ordinary differential equations with elliptic coefficients. They proved that all elliptic KdV algebro-geometry potentials are equivalent to the Picard potentials [21]. As examples they gave several elliptic potentials corresponding to genus $g=4,5$ ([21], Eqs.(5.40) and (5.43); also [20], Eqs.(3.86)-(3.89)). It is also worth mentioning that Krichever et.al. investigated the case of matrix KP equation [29] and the analogue research of [58] for matrix case was developed by Treibich in [60].
In addition, Belokolos and Enol’skii considered reduction from the general finite-gap solution to the DTV potential [6]. Later, they [7] described a technique on how elliptic finite-gap potentials are derived from the Matveev-Its formula for the finite-gap potential. Besides, Enol’skii and Kostov [11] discussed the conditions on the moduli of the corresponding algebraic curve under which the general finite-gap solution reduces to elliptic solitons. They indicated a completely algorithmic construction of the elliptic soliton moduli. There are also studies on the elliptic potentials related to the Ablowitz-Kaup-Newell-Segur spectral problem, for example, [52] from Smirnov and [22] by Gesztesy and Weikard.
The solutions we derived in this paper can be considered as special excitations of elliptic potentials without increasing genus. In illustration they are solitons living on an elliptic background. Strictly speaking, these solutions are not elliptic in $x$ any longer. We follow [38] and still employ the name “elliptic solitons” in this paper.